Cheryl climbed a set of stairs and stopped at the middle step. She then walked down 2 steps, up 4 steps, down 3 steps, and up 5 steps, and she was at the top of the stairs. How many steps are in the set of stairs?
Answer by macston(5194)

.
middle: m
down 2 steps: m-2
up 4 steps: (m-2)+4=m+2
down 3 steps: (m+2)-3=m-1
up 5 steps: (m-1)+5=m+4
.
Since there are 4 steps from middle to top.
there are 4 steps from middle to bottom,
for a total of 9 steps.
.
ANSWER: There are 9 steps.